OpenStudy (anonymous):

Solve the following using Cramer's Rule: 3y = 6x + 12 2y = 4x - 2

OpenStudy (turingtest):

what level math is this so I know how to explain this properly?

OpenStudy (anonymous):

99 smartness :O

OpenStudy (anonymous):

i am unable to type the solution by cramers rule.. hahahahah

OpenStudy (fools101):

here try this Web site if not let me know! http://www.mathway.com/problem.aspx?p=algebra ~Jess~

OpenStudy (turingtest):

@bipulrman plz don't provide the4 answers if you know cramers rule explain it, but don't just post answers as per the Code of Condduct

OpenStudy (anonymous):

It's Algebra! @TuringTest

OpenStudy (turingtest):

if it's basic algebra that makes it a pain to describe do you know how to find determinants?

OpenStudy (fools101):

1... 3y = 6x + 12 y=2x+4 2. 2y = 4x - 2 y=2x-1 I think this will be helpful to u if not let me know!! = ) ~Jess~

OpenStudy (anonymous):

I'm a bit rusty in math. Haven't been in school so I'm way behind. So no -__-

OpenStudy (anonymous):

@Fools101 thankyou !!

OpenStudy (fools101):

^-^ ur welcome!!!

OpenStudy (turingtest):

1) get all the variables on the same side of the equation please who me what you have after that

OpenStudy (turingtest):

show*

OpenStudy (anonymous):

As in get rid of the 3 and the 2 next to Y?

OpenStudy (turingtest):

not quite like for the first one I mean make it -6x+3y=12 see what I mean (x's and y's on the same side)? what would the second equation be?

OpenStudy (turingtest):

honestly, math is not one of those things you can just cram on at the last minute if you are "way behind" Cramer's rule is going to be very difficult to describe, sorry...

OpenStudy (anonymous):

-4x - 2y =2

OpenStudy (anonymous):

& I'm not 'cramming'. I'm asking for help. Trying to understand more then what the online slides are showing.

OpenStudy (turingtest):

well that is correct, so nicely done the next step is to write the coefficients as what's called an "augmented matrix" which I will draw....

OpenStudy (turingtest):

here is our augmented matrix: